OpenStudy (anonymous):

Using the formula P = 2(a + b), find a when P = 25 and b = 9.

OpenStudy (zepp):

Simply plug in the value and solve.

OpenStudy (zepp):

And don't expect an answer.

OpenStudy (jiteshmeghwal9):

p=2(a+b) put he given values in the equation 25=2(a+9) now solve it:)

OpenStudy (zepp):

@SmartWish @pinkpassionn Let him answer.

OpenStudy (anonymous):

okay @zepp .. just plug in the values like zepp said :) @beezz12

OpenStudy (anonymous):

3.5

OpenStudy (anonymous):

what did you do to get 3.5 ? @beezz12

OpenStudy (anonymous):

3.5+9=12.5 12.5*2=25

OpenStudy (zepp):

25=2(a+9) 25=2a+18 25-18=2a You should have done that, I don't know where your 2.5 come from

OpenStudy (anonymous):

that answer wasnt a choice, so I got 3.5

OpenStudy (zepp):

3.5 is right

OpenStudy (anonymous):

: ) good .